OpenStudy (anonymous):

Ill give you a medal ! Which is a translation of the sentence into an equation? 1. The sum of a number and six is seven times the number. A: n = 7n + 6 B: 7(n + 6) = n C: n − 6 = 7n D: n + 6 = 7n

OpenStudy (haleyelizabeth2017):

n+6=7n

OpenStudy (anonymous):

2. Which value for the number makes this statement true? The quotient of a number and five is eight. A: 4 B: 8 C: 40 D: 80

OpenStudy (haleyelizabeth2017):

40

OpenStudy (anonymous):

3. (last one) Tran drove his car at a constant speed of 55 miles per hour. Which is an equation for the distance Tran drove? A: 55d = t B: d= 55t C: t/d = 55 D: 55/d = t

OpenStudy (haleyelizabeth2017):

I don't know this one. If I were to guess, I would choose D.....

OpenStudy (anonymous):

Thank you! You've got +1 fan :) @haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

:) np